8. The distance (in km) of 40 engineers from their residence to their place of work were found as follows:

6 3 10 20 26 11 13 7 12 31
19 10 12 17 18 11 32 17 16 2
7 9 7 8 3 6 12 16 18 3
12 14 2 9 6 15 16 7 6 12
What is the empirical probability that an engineer lives:
(i) less than 7 km from her place of work?
(ii) more than or equal to 7 km from her place of work?
(iii,) within ikm from her place of work?

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Total number of engineers = 40. (i) Number of engineers having distance of less than 7 km from their residence to the place of work = 9. PROBABILITY 237 .. Probability that an engineer lives less than 7 km from her place of work = 9 . 40 (ii') Number of engineers having distance of more than or equal to 7 km from their residence to the place of work = 40 - 9 = 31. :. Probability that an engineer lives more than or equal to 7 km from her place of work = 31 40 (iii) There is no engineer within -i km from her place of work. :. Probability that an engineer lives within -ikm from her place of work = 0.